OpenStudy (anonymous):

marry bought 2 1/2 yards of fabric for 12.50 per yard. she also baught 2 1/2 yards of silk fabric for 30$ per yard. how much did mary spend on fabric all together?

OpenStudy (anonymous):

i change it to an improper fraction then multiply that by 12.50 right?

OpenStudy (anonymous):

price * unit is your answer 2 1/2 = 5/2 , 5/2 * 12.5 + 5/2 * 30 = 5/2 * ( 12.5 + 30 )

OpenStudy (anonymous):

106.25 thank you soo much!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

106.25 thank you soo much!